From GlueXWiki
Jump to: navigation, search

January 30, 2014 FDC meeting


  1. Installation [1],[2] (all)
  2. Engineering December 2013 survey
  3. Electronics (Chris, Nick)
  4. Mini-DAQ results FDC E-log (Beni)
  5. Other


Participants: Fernando, Dave, Nick, Chris, Simon, Beni, Eugene, and Lubomir.


- Casey, Tina and Anatoly finished cabling the FDC at the detector side and it looks nice (first picture attached). On Tuesday the platform on the right side was opened and today Nick and Chris started routing the right bottom quadrant to the electronics. They are done with package #4 working now on #3.

- HV: was stable since the beginning of the week at +2200V/-500V, except for 6 channels at +2200/-450V and three channels at +2200/-500V, out of 96 total number of HV channels. Such change of the field voltage on these 9 channels affects the biggest drift times/distances where the resolution is in any case poor, so the average track resolution will not be affected. Note also that +2200V is the upper limit of the HV range and we may not need to be there.

- Gas: since Tuesday we are running gas at 400ccpm (according to the mass flow controller set point) compared to 800ccpm before, because we don't have more bottles and the next ones are expected to come next week. This morning somebody working at upstream face of the magnet, pulled the exhaust gas line and we found (also in the morning) the bubbler and a "quick-connect" connector disconnected. As a result the pressure went down to ~20Pa from ~60Pa. The pressure recovered in about two hours. At the end of last week one of the bubblers stopped bubbling and the pressure started increasing (as usual), however, after inserting a tube at the output of the bubbler going inside the oil (second picture attached) it started bubbling again with high rate and the bubbling is very stable since then.

- Dave proposed to put a plastic shield at the downstream side of the FDC to protect the gas and other FDC stuff there. Also expect that the test of the BCAL water cooling system which is next to the FDC, will start next week (need an umbrella!).


Bill was busy at another meeting, but he communicated with Lubomir:

- We have now the report (attached) of the survey of the 16 reflectors done in December when the FDC was inside the magnet. In z-direction, the deviations from the nominal are not more than ~0.5 mm resulting of a pitch and yaw less than 0.4 mrad. Such deviations will have just a little contribution to the resolution even if not corrected for. It would be interesting to compare with the survey results when the FDC was outside of the magnet.

- Bill received an answer from BROOKS: the mass flow controllers are to be blamed and now both the set and the read-out values are different from the actual flow. Beni: all the mass flow controllers that have been re-calibrated in the second batch, 7 in total, have the same problem. We have to send them for fixing a.s.a.p., still we want to keep one to be used in the temporary gas system that we have now for the FDC. The two mass flow controllers that have been re-calibrated first, are working fine (now in EEL126).

- Bill found a company (Hexagon Metrology) in Charlotte, NC that has the machine (MCC) to scan the cathode. Bill is arranging a trip to there with the cathode to do the measurement in the next few weeks.


- The adapter (a PCB) from the 3M connector (that is on the signal cables) to the differential probe is ready. We started immediately using it for testing the cables with a scope, instead of the shaper board that sees only 16 channels and has additional shaping. Fernando also wanted to look with the scope at the FDC signals next week.

- For the next meeting Lubomir will make a list of all the problematic cables/connectors/pre-amps as identified so far with the scope testing or with the mini-DAQ.

- Since we will have all the cables routed to the right racks we will start at the same time plugging the cables to the modules. Not all the modules are in place, but Nick and Chris will install the missing modules (now Cody has them) and Fernando agreed.

- The grounding cable is ready, needs to be attached to the FDC.

- Dave will work on the hardware shutdown of the low voltage crate (when the cooling fails), with a help from Fernando to build a simple PCB for that.


- Beni presented his latest results from the mini-DAQ (E-log entries 131-133) for the last quadrant (bottom left) the first two packages. TDC results: there are several missing wire channels in one card. fADC results: there's one very noisy card and several missing channels. Lubomir and Anatoly looked at the problematic channels with a scope. The "missing wire" channels were fixed by re-seating the pre-amp and bending the cooling tube so that it doesn't apply pressure on the card. Fernando and Chris discussed that such pressure is unlikely to damage the connector on the wire PCB, but rather it can damage the connector on the pre-amp. The noisy cathode card indeed has a low amplitude (10-15mV) few hundred Hz noise. This is a known problem at the place where the HV via faces the strip connector. The missing strip channels were either due to a bad cable/connectors or, in two cases, the signals are smaller due to the conductive tape problem.

- Fernando: Cody is very close to be able to download the fADC125 firmware through the VME, the fast readout is ready after that, in February, the full crate tests will be performed. We expect the FDC to be cabled completely by the end of the next week, so if there are no issues with the firmware we can start the mini-DAQ testing.